OpenStudy (anonymous):

The formula for the volume (V) of a box is V = lwh. Find h, if V = 0.0144 m3, l = 0.08 m, and w = 0.6 m.

OpenStudy (anonymous):

w = 3 m , i got that, but it was rong

OpenStudy (mayankdevnani):

0.144=0.08*0.6*h 0.144=0.048*h h=0.144/0.048=0.3m

OpenStudy (mayankdevnani):

@sbanuelos ok